Suppose the length of a clock's pendulum is decreased by 1.100%, exactly at noon one day. What time will it read 24.00 hours later, assuming it the pendulum has kept perfect time before the change? Perform the calculation to at least five-digit precision. Enter you time as hh:mm:ss AM/PM. Thank you.
Answer #1

here,

the new length of pendulum , l = (100 - 1.1)/100 * l0

l = 0.989 * l0

the initial time period , T0 = 24 hrs

let the new time time period be T

T /T0 = sqrt(l/g) /(sqrt(l0/g))

T /24 = sqrt(0.989)

T = 23.867 h = 23 h 52 min 3 s

the new time is (11:52:3) AM